OpenStudy (anonymous):

Find the intersection point(s) of the graphs of: y=x^2 and -6+y=-1x Plot the intersection points (not the graphs of the equations) on the axes below.

OpenStudy (anonymous):

Well, to find intersection points, just solve for y and set them equal to each other, then solve for x y=x^2 y=-1x+6 x^2=-1x+6 x^2+1x-6 = 0

OpenStudy (anonymous):

-x+6=x^2 x^2+x-6=0 (x+3)(x-2)=0 They intersect with x = -3 and at x = 2 Intersection at the points (-3, 9) and (2, 4)

OpenStudy (anonymous):

TBates: Where did you get 9 and 4 from?

OpenStudy (anonymous):

From substituting in x in the original equations (-3)^2 = 9 (2)^2 = 4

OpenStudy (anonymous):

I'm graphing it wrong for some reason on my online homework. Can you draw it out please?

OpenStudy (anonymous):

It seems to just want the points plotted on the graph, so it's just a point at (-3, 9) and (2, 4)

OpenStudy (anonymous):